a certain medicine is prescribed in an amount proportional to a patient's body wieght. if a patient weighing 71 kilograms requires 213 milligrams of this medicine, then amount of medicine required for a patient weighing 78 kilograms is:
Answer by stanbon(48568)

if a patient weighing 71 kilograms requires 213 milligrams of this medicine, then amount of medicine required for a patient weighing 78 kilograms is:
----------------
x/78 = 213/71
x= 78(213/71)
x=234 milligrams
----------
Cheers,
Stan H.